Find the equation of the cyclinder with base x²+y²+z²+3x+3y-z=1, x+y+2z=2.
1
Expert's answer
2019-07-16T08:47:55-0400

Given plane equation is (x+y+2z-2)=0, sphere equation is (x²+y²+z²+3x+3y-z)=1


"x^2+2\cdot{3 \over 2}x+{9 \over 4}-{9 \over 4}+y^2+2\cdot{3 \over 2}y+{9 \over 4}-{9 \over 4}+"


"+z^2-2\cdot{1 \over 2}z+{1 \over 4}-{1 \over 4}=1"

"\big(x+{3 \over 2}\big)^2+\big(y+{3 \over 2}\big)^2+\big(z-{1 \over 2}\big)^2={23 \over 4}"

"Centre (-u,-v,-w)=\big(-{3 \over 2},-{3 \over 2},{1 \over 2}\big)"


"Radius\ \ r=\sqrt{{23 \over 4}}={\sqrt{23} \over 2}"


Length of the perpendicular from "\big(-\dfrac{3}{2},-\dfrac{3}{2},\dfrac{1}{2}\big)" to the plane (x+y+2z-2) =0 is 


"OM=\dfrac{\big|ax_1+by_1+cz_1+d\big|}{\sqrt{a^2+b^2+c^2}}"

Here, "a=1,b=1,c=2,d=-2,x_1=-\dfrac{3}{2},y_1=-\dfrac{3}{2},z_1=\dfrac{1}{2}"



"OM=\dfrac{\big|-\dfrac{3}{2}-\dfrac{3}{2}+2\cdot\dfrac{1}{2}-2\big|}{\sqrt{1^2+1^2+2^2}}=\dfrac{2\sqrt6}{3}"


"AM^2=OA^2-OM^2=({\sqrt{23} \over 2})^2-({2\sqrt6 \over 3})^2={37 \over 12}"

The equation of the axis of the cylinder is 


"{x-\alpha \over l}={y-\beta \over m}={z-\gamma \over n}"

Here, "\alpha=-\dfrac{3}{2},\beta=-\dfrac{3}{2},\gamma=\dfrac{1}{2}, l=1,m=1,n=2."


"{x+\dfrac{3}{2} \over 1}={y+\dfrac{3}{2} \over 1}={z-\dfrac{1}{2} \over 2}"

"l^2+m^2+n^2=6"

"Radius\ \ r^*=AM=\sqrt{{37 \over 12}}={\sqrt{111} \over 6}"

The equation of Right Circular Cylinder is, 


"[n(y-\beta)-m(z-\gamma)]^2+[l(z-\gamma)-n(x-\alpha)]^2+""+[m(x-\alpha)-l(y-\beta)]^2=(r^*)^2(l^2+m^2+n^2)"

"[2(y+{3 \over 2})-1(z-{1\over 2})]^2+[1(z-{1 \over 2})-2(x+{3\over 2})]^2+""+[1(x+{3 \over 2})-1(y+{3 \over 2})]^2=({\sqrt{111} \over 6})^2(6)"


"4y^2+z^2+{49 \over 4}-4yz+14y-7z+z^2+4x^2+{49 \over 4}-""-4xz-7z+14x+x^2-2xy+y^2={37 \over 2}"



"5x^2+5y^2+2z^2-2xy-4xz-4yz+14x+14y-14z+6=0"